OpenStudy (anonymous):

Evaluate the expression 7x exonent 2y, when x = 5and y= 3. 210 315 525 728

Mehek (mehek14):

is it supposed to be \(7x^{2y}?\)

OpenStudy (anonymous):

the y is not the expo nent tho

OpenStudy (anonymous):

exonent

Mehek (mehek14):

\(7x^2y\)?

OpenStudy (anonymous):

yes

Mehek (mehek14):

ok so \(7x^2y=7*x^2*y\) you are given x = 5 and y = 3 plug those in \(7*5^2*3\)

OpenStudy (anonymous):

525?

Mehek (mehek14):

correct

OpenStudy (anonymous):

ok thxs

Mehek (mehek14):

yw :)